# PUL-SLAM: Path-Uncertainty Co-Optimization with Lightweight Stagnation Detection for Efficient Robotic Exploration

### v1

@@ Abstract
Existing Active SLAM methodologies face issues such as slow exploration speed and suboptimal paths. To address these limitations, we propose a hybrid framework combining a Path-Uncertainty Co-Optimization Deep Reinforcement Learning framework and a Lightweight Stagnation Detection mechanism. The Path-Uncertainty Co-Optimization framework jointly optimizes travel distance and map uncertainty through a dual-objective reward function, balancing exploration and exploitation. The Lightweight Stagnation Detection reduces redundant exploration through Lidar Static Anomaly Detection and Map Update Stagnation Detection, terminating episodes on low expansion rates. Experimental results show that compared with the frontier-based method and RRT method, our approach shortens exploration time by up to 65 and reduces path distance by up to 42, significantly improving exploration efficiency in complex environments while maintaining reliable map completeness. Ablation studies confirm that the collaborative mechanism accelerates training convergence. Empirical validation on a physical robotic platform demonstrates the algorithm’s practical applicability and its successful transferability from simulation to real-world environments.
1 Introduction
@@
Active Simultaneous Localization and Mapping (Active SLAM) requires robots to simultaneously perform three critical tasks in unknown environments: environmental mapping, self-localization, and exploration path planning, thereby enabling efficient environmental exploration[1]. This technology has proven indispensable in critical application scenarios such as disaster rescue[2], planetary exploration[3], underground mine exploration[4, 5], and infrastructure inspection[6], particularly in environments that are inaccessible or hazardous to humans, where fully autonomous robotic exploration systems can substantially improve task execution efficiency while minimizing personnel risks[7]. The fundamental challenge of Active SLAM lies in achieving a dynamic equilibrium between exploration (discovering new areas) and exploitation (revisiting known regions to reduce localization and mapping uncertainty): robots must rapidly discover new areas while simultaneously conducting sufficient exploration of known regions to minimize uncertainties in positioning and mapping. Unlike traditional SLAM approaches that focus exclusively on mapping and localization accuracy, Active SLAM introduces significant additional complexity through path planning and exploration strategy decision-making, thereby transforming the problem into a highly complex multi-objective optimization challenge.
However, existing methods still face several issues, such as slow exploration speed with significant time consumption during path planning, suboptimal and circuitous exploration paths, as well as limited scene adaptability and scalability leading to increased computational costs in complex environments.
Active SLAM methodologies exhibit significant diversity in exploration strategies, yet each faces distinct limitations in practical deployment. Frontier-based exploration-pioneered by Yamauchi[8] and enhanced through techniques like wavefront frontier detection (WFD)[9]—remains widely used to this day and can be found in numerous recent works[10, 11]. However, this approach incurs high computational overhead in large-scale environments, and lacks global trajectory optimization mechanisms. Random sampling-based methods, such as Rapidly-exploring Random Trees (RRT)[12] and Probabilistic Roadmaps (PRM)[13], prioritize viewpoint diversity through stochastic sampling but frequently produce tortuous, energy-inefficient paths, exhibit incomplete area coverage due to undersampling. Information-theoretic approaches—exemplified by Shannon entropy maximization[14] and mutual information[15] optimization—systematically reduce map uncertainty through probabilistic modeling. Nevertheless, real-time computation of information gain metrics remains computationally prohibitive, performance degrades under sensor noise , and path efficiency is often sacrificed for uncertainty reduction. Notably, although deep reinforcement learning (DRL) has introduced novel paradigms for end-to-end decision learning[16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], its application in Active SLAM continues to face significant challenges related to reward sparsity and convergence difficulties. Traditional DRL approaches often overly emphasize single objectives (e.g., map coverage ratio or uncertainty reduction) while neglecting the critical trade-offs between multiple competing objectives, which prevents the achievement of optimal overall performance in complex environments. Furthermore, existing systems lack dynamic detection and correction mechanisms for abnormal exploration behaviors, causing robots to become trapped in inefficient exploration states when encountering complex obstacles, thereby hindering their ability to adaptively adjust exploration strategies.
@@
To address these limitations, we propose an innovative dual-layer collaborative optimization framework that jointly models path optimization and uncertainty reduction, while introducing a lightweight stagnation detection mechanism to enhance the system’s adaptability in complex environments. The main contributions of this study include:
@@
Path-Uncertainty Co-Optimization DRL Framework: We propose a novel deep reinforcement learning framework that jointly optimizes travel distance and map uncertainty through a dual-objective reward function, balancing exploration and exploitation.
@@
Experimental results show that compared with the bounded-based method and RRT method, the time is shortened by up to 60%, and the path is shortened by up to 40%, which significantly improves the exploration efficiency in complex environments.
@@ 2 RELATED WORK
In the field of robotic autonomous exploration, the design of reward functions and termination conditions constitutes the core decision-making mechanism of Active SLAM systems, directly influencing exploration efficiency and map quality. The following sections systematically review the research progress in these two critical aspects.
2.1 Exploration Strategy Reward Function
@@
2.1.1 Map-Completeness-Based Reward Mechanisms
@@
This approach motivates robots to achieve comprehensive environmental traversal through coverage increment incentives. For instance, Zhao et al.[18] decomposed rewards into map completeness, exploration rewards, and exploitation rewards to holistically incentivize exploration behaviors. Chaplot et al.[25] directly designed reward functions based on increases in covered area. These methods offer intuitive interpretability and ensure systematic environmental traversal. However, they often lead to suboptimal path planning in complex environments, particularly in obstacle-dense regions where robots may become trapped in inefficient repetitive exploration due to excessive focus on local coverage.
2.1.2 Environment-Uncertainty-Based Reward Mechanisms
@@
These methods leverage information entropy reduction or feature metrics of SLAM covariance matrices to drive active exploration. Chen et al.[20] proposed a reward function integrating map information gain, control rewards, exploration completion rewards, and collision penalties. Alcalde et al.[21] and A. Placed et al.[23] adopted the D-optimality criterion to quantify localization and mapping uncertainties, embedding this metric into reward design. These approaches possess solid theoretical foundations and effectively reduce map uncertainty while improving localization accuracy. However, sensor noise can degrade performance, and path efficiency is often sacrificed to reduce uncertainty, resulting in excessive detours during exploration.
2.1.3 Other Reward Mechanisms
@@
Beyond the two primary mechanisms, several innovative reward designs have been proposed. For instance, Cao et al.[19] designed a composite reward function incorporating frontier point counts, path length penalties, and task completion incentives. In a related approach, Botteghi et al.[22] introduced an intrinsic curiosity-driven mechanism to encourage exploratory behavior. Similarly, Zhu et al.[24] employed negative penalties proportional to path length to promote shorter, more efficient trajectories.
@@
Notably, path length, as a critical metric of exploration efficiency, has been rarely systematically incorporated into reward function design in existing literature. While Cao et al.[19] and Zhu et al.[24] introduced path length penalty terms in their respective works, no prior studies have proposed jointly optimizing path length and map uncertainty as a dual-objective framework. The path-uncertainty co-optimization framework proposed in this study integrates both metrics into a unified reward function, dynamically balancing the trade-off between exploration and exploitation. This approach effectively addresses the suboptimal path planning issues resulting from existing methods’ excessive focus on single objectives such as coverage area or map entropy, thereby providing a more comprehensive and effective decision-making mechanism for autonomous robotic exploration.
2.2 Task Termination Mechanism
@@
2.2.1 Environment-Triggered Termination Mechanisms
Collision detection represents a typical example of this category, where tasks are terminated when the robot-obstacle distance falls below a predefined threshold (e.g., 0.2 meters) [18, 20, 22, 21, 23]. This approach effectively prevents robots from continuing operation in hazardous environments, ensuring system safety. However, this mechanism lacks dynamic awareness of exploration progress, making it difficult to adapt to environments of varying complexity.
2.2.2 Task-Driven Termination Mechanisms
Exploration completion thresholds (e.g., coverage ratio ) have been validated and applied across multiple studies[18, 19, 20, 22, 23, 24, 25]. This mechanism ensures exploration tasks reach predefined objectives, but the threshold settings lack adaptability, making them unsuitable for environments of different scales and complexities. In simple environments, termination may occur too early; in complex environments, the threshold may never be reached, potentially resulting in indefinite task duration. Furthermore, this mechanism lacks dynamic awareness of exploration progress, making it difficult to adapt to environments of varying complexity.
2.2.3 Resource-Constrained Termination Mechanisms
Fixed step limits or time ceilings ensure computational efficiency, as implemented in[20, 21, 25]. These methods prevent indefinite exploration through predefined resource constraints but lack dynamic awareness of exploration progress. This can result in critical regions remaining unexplored before resource exhaustion or premature termination when resources are still available.
@@ 3.1.1 Reward Function
where is a task-dependent scale factor, is the D-optimality criterion[21], represents the newly added map area at time , and is the path penalty term, which is defined as follows:
@@ 3.1.2 Observation Space
At time step , we uniformly sample laser measurements to obtain ranging values, as illustrated in Fig. 1, yielding range values normalized to . By reducing the number of sampling points , computational complexity is effectively reduced while maintaining sufficient environmental representational capacity, allowing the algorithm to run in real-time on resource-constrained mobile robot platforms.
The observation space for the reinforcement learning agent consists of two components: the normalized laser scan vector and the current map coverage ratio . This design choice integrates local perception with global state awareness: the laser scan vector captures fine-grained geometric details of the immediate surroundings, furnishing the agent with real-time sensory input for decision-making, while the cumulative map coverage ratio serves as a global indicator of exploration progress, thereby enabling the agent to maintain a coherent understanding of its spatial context.
@@ 3.1.3 Action Space
We adopt the Proximal Policy Optimization (PPO)[27] algorithm for policy learning. The observation space is structured as a dictionary comprising 24-dimensional normalized LiDAR readings and a 1-dimensional map coverage metric, resulting in a 25-dimensional joint input vector. This concatenated observation is processed by a shared backbone network consisting of two fully connected layers, each with 64 neurons and Tanh activation functions. The shared representation is then fed into two separate heads: a policy head that outputs logits over three discrete actions (forward, turn left, turn right), and a value head that estimates the scalar state-value function. All network parameters are jointly optimized in an end-to-end manner via gradient-based updates.
@@ 3.2.1 Lidar Static Anomaly Detection
Robotic exploration can exhibit intentional pausing (e.g., stationary behavior to maximize reward in RL) or motion failure (e.g., collisions or wheel slippage). Undetected motion failures degrade learning efficiency by slowing convergence and reinforcing suboptimal policies. To address this, we propose Lidar Static Anomaly Detection: a lightweight method that identifies motion failures via cosine similarity between consecutive LiDAR scans. Unlike odometry/IMU-based approaches, this method uses raw environmental perception data, maintaining reliability during motor idling or wheel slippage. By operating on normalized scan vectors, it inherently rejects localized environmental changes. LSAD achieves real-time stagnation detection, with negligible impact on exploration performance.

### v2

@@ Abstract
Existing Active SLAM methodologies face issues such as slow exploration speed and suboptimal paths. To address these limitations, we propose a hybrid framework combining a Path-Uncertainty Co-Optimization Deep Reinforcement Learning framework and a Lightweight Stagnation Detection mechanism. The Path-Uncertainty Co-Optimization framework jointly optimizes path length and pose uncertainty through a dual-objective reward function, balancing exploration and exploitation. The Lightweight Stagnation Detection reduces redundant exploration through Lidar Static Anomaly Detection and Map Update Stagnation Detection, terminating episodes on low expansion rates. Experimental results show that compared with the frontier-based method and RRT method, our approach shortens exploration time by up to 65 and reduces path distance by up to 42, significantly improving exploration efficiency in complex environments while maintaining reliable map completeness. Ablation studies confirm that the collaborative mechanism accelerates training convergence. Empirical validation on a physical robotic platform demonstrates the algorithm’s practical applicability and its successful transferability from simulation to real-world environments.
keywords:
Active SLAM, Deep Reinforcement Learning, Autonomous Exploration, Intelligent Robotics1 Introduction
@@
Active Simultaneous Localization and Mapping (Active SLAM) requires robots to simultaneously perform three critical tasks in unknown environments: environmental mapping, self-localization, and exploration path planning, thereby enabling efficient environmental exploration[placed2023survey]. This technology has proven indispensable in critical application scenarios such as disaster rescue[niroui2019deep], planetary exploration[nahavandi2025comprehensive], underground mine exploration[wang2025multi, ackerman2022robots], and infrastructure inspection[zhou2023racer], particularly in environments that are inaccessible or hazardous to humans, where fully autonomous robotic exploration systems can substantially improve task execution efficiency while minimizing personnel risks[azpurua2023survey]. The fundamental challenge of Active SLAM lies in achieving a dynamic equilibrium between exploration (discovering new areas) and exploitation (revisiting known regions to reduce localization and mapping uncertainty): robots must rapidly discover new areas while simultaneously conducting sufficient exploration of known regions to minimize uncertainties in positioning and mapping. Unlike traditional SLAM approaches that focus exclusively on mapping and localization accuracy, Active SLAM introduces significant additional complexity through path planning and exploration strategy decision-making, thereby transforming the problem into a highly complex multi-objective optimization challenge.
Active SLAM methodologies exhibit significant diversity in exploration strategies. Traditional frontier-based methods [yamauchi1997frontier, keidar2014efficient, wang2024exploration, saravanan2024fit], random sampling techniques like RRT [umari2017autonomous, faust2018prm, wu2019autonomous], and information-theoretic approaches [whaite1997autonomous, asgharivaskasi2023semantic, stachniss2005information, carrillo2012comparison] have long dominated exploration strategies in unknown environments. These methods feature intuitive implementation architectures and are relatively easy to deploy. However, they employ static decision-making strategies that cannot adapt to dynamic environmental factors such as environmental complexity and exploration progress, resulting in unstable performance in varied environments. Furthermore, they typically optimize only a single objective (frontier-based methods focus on coverage ratio, RRT emphasizes path feasibility, and information-theoretic approaches prioritize uncertainty reduction), while neglecting the multi-objective trade-offs involving exploration efficiency, energy consumption, and other factors. This limitation, combined with their lack of global perspective, often leads to redundant backtracking paths in cluttered spaces. Recent deep reinforcement learning (DRL) methods [zhao2024autonomous, zhou2024indoor, zhao2024exploration, cao2023ariadne, chen2024lidar, alcalde2022slam, botteghi2021curiosity, placed2020deep, zhu2018deep, chaplot2020learning, zhao2025multirobot] have demonstrated considerable potential in autonomous exploration domains. These approaches enable agents to interact with the environment and continuously refine their exploration strategies based on reward feedback, thereby optimizing the decision-making process. However, the performance of these methods is highly dependent on the design of the reward function. A well-designed reward function can steer the agent toward an optimal policy. Conversely, an improper design may result in the agent learning suboptimal strategies that significantly diverge from efficient exploration paths.
@@
Despite these diverse paradigms, current systems still struggle to achieve an optimal balance between exploration efficiency and mapping accuracy, often suffering from slow exploration speeds and suboptimal trajectories. Additionally, existing systems lack mechanisms to dynamically detect and correct inefficient exploration behaviors, causing robots to become trapped in local oscillations or redundant exploration patterns when encountering complex obstacles. To address these limitations, we propose an innovative dual-layer collaborative optimization framework that jointly models path optimization and uncertainty reduction, while introducing a lightweight stagnation detection mechanism to enhance the system’s adaptability in complex environments. The main contributions of this study include:
@@
Path-Uncertainty Co-Optimization DRL Framework: We propose a novel deep reinforcement learning framework that jointly optimizes travel distance and pose uncertainty through a dual-objective reward function, balancing exploration and exploitation.
@@
Extensive simulation and real-world experiments: Experimental results show that compared with the frontier-based method and RRT method, the time is shortened by up to 65%, and the path is shortened by up to 42%, which significantly improves the exploration efficiency.
@@ 2 RELATED WORK
2.1 Traditional Active SLAM
Traditional Active SLAM methodologies can be broadly categorized into frontier-based exploration, random sampling techniques, and information-theoretic approaches, each with distinct characteristics and limitations that inform our current research direction.
Frontier-based exploration, pioneered by Yamauchi [yamauchi1997frontier], represents one of the most influential paradigms in robotic exploration. This approach identifies boundary regions between known and unknown areas (frontiers) and directs robots toward these locations to maximize information gain. Enhanced variants like Wavefront Frontier Detection (WFD) [keidar2014efficient] improved computational efficiency but still suffer from significant drawbacks in complex environments: the algorithm tends to generate suboptimal paths with excessive backtracking, struggles with large-scale environments due to growing computational overhead, and lacks mechanisms for global trajectory optimization. Despite these limitations, frontier-based methods remain widely adopted in recent works [wang2024exploration, saravanan2024fit] due to their intuitive implementation and reliable coverage performance.
Random sampling-based methods, particularly Rapidly-exploring Random Trees (RRT) [umari2017autonomous], offer an alternative exploration strategy by constructing search trees through random sampling of the configuration space. While RRT variants excel at finding feasible paths in high-dimensional spaces and complex obstacle arrangements, they exhibit critical shortcomings for exploration tasks: the resulting trajectories are often tortuous and energy-inefficient, coverage completeness is compromised due to undersampling in narrow passages, and the stochastic nature of sampling leads to inconsistent exploration patterns. These limitations become particularly pronounced in cluttered indoor environments where systematic coverage is essential.
Information-theoretic approaches employ metrics like Shannon entropy [asgharivaskasi2023semantic] and mutual information [stachniss2005information] to quantify and reduce mapping uncertainty through probabilistic modeling. Though theoretically sound, these methods face practical challenges including prohibitive computational costs for real-time information gain calculation, sensitivity to sensor noise, and frequent sacrifice of path efficiency for uncertainty reduction.
These traditional approaches establish important foundations but reveal significant gaps in handling the multi-objective nature of exploration. Their limitations in path optimality, computational efficiency, and adaptive decision-making motivate our development of a hybrid framework that preserves the strengths of systematic exploration while addressing these fundamental challenges through modern learning techniques.
2.2 DRL-based Active SLAM
In the field of DRL-based robotic autonomous exploration, the design of reward functions and termination conditions constitutes the core decision-making mechanism of Active SLAM systems, directly influencing exploration efficiency and map quality. The following sections systematically review the research progress in these two critical aspects.
2.2.1 Exploration Strategy Reward Function
@@
Map-Completeness-Based Reward Mechanisms
@@
These methods motivate robots to achieve comprehensive environmental traversal through coverage increment incentives. For instance, Zhao et al.[zhao2024exploration] decomposed rewards into map completeness, exploration rewards, and exploitation rewards to holistically incentivize exploration behaviors. Chaplot et al.[chaplot2020learning] directly designed reward functions based on increases in covered area. These methods offer intuitive interpretability and ensure systematic environmental traversal. However, they often lead to suboptimal path planning in complex environments, particularly in obstacle-dense regions where robots may become trapped in inefficient repetitive exploration due to excessive focus on local coverage.
Environment-Uncertainty-Based Reward Mechanisms
@@
These methods leverage information entropy reduction or feature metrics of SLAM covariance matrices to drive active exploration. Chen et al.[chen2024lidar] proposed a reward function integrating map information gain, control rewards, exploration completion rewards, and collision penalties. Alcalde et al.[alcalde2022slam] and A. Placed et al.[placed2020deep] adopted the D-optimality criterion to quantify localization and mapping uncertainties, embedding this metric into reward design. These approaches possess solid theoretical foundations and effectively reduce map uncertainty while improving localization accuracy. However, sensor noise can degrade performance, and path efficiency is often sacrificed to reduce uncertainty, resulting in excessive detours during exploration.
Other Reward Mechanisms
@@
Beyond the two primary mechanisms, several innovative reward designs have been proposed. For instance, Cao et al.[cao2023ariadne] designed a composite reward function incorporating frontier point counts, path length penalties, and task completion incentives. In a related approach, Botteghi et al.[botteghi2021curiosity] introduced an intrinsic curiosity-driven mechanism to encourage exploratory behavior. Similarly, Zhu et al.[zhu2018deep] employed negative penalties proportional to path length to promote shorter, more efficient trajectories.
@@
Notably, path length, as a critical metric of exploration efficiency, has been rarely systematically incorporated into reward function design in existing literature. While Cao et al.[cao2023ariadne] and Zhu et al.[zhu2018deep] introduced path length penalty terms in their respective works, no prior studies have proposed jointly optimizing path length and pose uncertainty as a dual-objective framework. The path-uncertainty co-optimization framework proposed in this study integrates both metrics into a unified reward function, dynamically balancing the trade-off between exploration and exploitation. This approach effectively addresses the suboptimal path planning issues resulting from existing methods’ excessive focus on single objectives such as coverage area or map entropy, thereby providing a more comprehensive and effective decision-making mechanism for autonomous robotic exploration.
2.2.2 Task Termination Mechanism
@@
Environment-Triggered Termination Mechanisms
Collision detection represents a typical example of this category, where tasks are terminated when the robot-obstacle distance falls below a predefined threshold (e.g., 0.2 meters) [zhao2024exploration, chen2024lidar, botteghi2021curiosity, alcalde2022slam, placed2020deep]. This approach effectively prevents robots from continuing operation in hazardous environments, ensuring system safety. However, this mechanism lacks dynamic awareness of exploration progress, making it difficult to adapt to environments of varying complexity.
Task-Driven Termination Mechanisms
Exploration completion thresholds (e.g., coverage ratio ) have been validated and applied across multiple studies[zhao2024exploration, cao2023ariadne, chen2024lidar, botteghi2021curiosity, placed2020deep, zhu2018deep, chaplot2020learning]. This mechanism ensures exploration tasks reach predefined objectives, but the threshold settings lack adaptability, making them unsuitable for environments of different scales and complexities. In simple environments, termination may occur too early; in complex environments, the threshold may never be reached, potentially resulting in indefinite task duration. Furthermore, this mechanism lacks dynamic awareness of exploration progress, making it difficult to adapt to environments of varying complexity.
Resource-Constrained Termination Mechanisms
Fixed step limits or time ceilings ensure computational efficiency, as implemented in[chen2024lidar, alcalde2022slam, chaplot2020learning]. These methods prevent indefinite exploration through predefined resource constraints but lack dynamic awareness of exploration progress. This can result in critical regions remaining unexplored before resource exhaustion or premature termination when resources are still available.
@@ 3 APPROACH
As illustrated in Fig. 1, the PUL-SLAM system features a dual-layer collaborative optimization architecture, integrating a high-level DRL-based decision-making framework with a low-level Lightweight Stagnation Detection mechanism. In the operational pipeline, the SLAM module first processes LiDAR inputs to generate a real-time occupancy grid map alongside robot pose estimation. Subsequently, the Lightweight Stagnation Detection module, which consists of LiDAR Static Anomaly Detection and Map Update Stagnation Detection, monitors for inefficient behaviors. The former identifies motion anomalies via consecutive frame similarity analysis, while the latter tracks map expansion rates. If any abnormal state is identified, the system promptly halts the current exploration episode and performs an environment reset to avoid reinforcing inefficient strategies. Simultaneously, the DRL framework leverages a path-uncertainty co-optimization strategy to dynamically trade off exploration efficiency with pose uncertainty, thereby producing optimal motion commands to drive the robot.
@@ 3.1.1 Reward Function
where is a task-dependent scale factor, is the D-optimality criterion[carrillo2012comparison, alcalde2022slam], represents the newly added map area at time , and is the path penalty term, which is defined as follows:
@@ 3.1.2 Observation Space
At time step , we uniformly sample laser measurements to obtain ranging values, as illustrated in Fig. 2, yielding range values normalized to:
By reducing the number of sampling points , computational complexity is effectively reduced while maintaining sufficient environmental representational capacity, allowing the algorithm to run in real-time on resource-constrained mobile robot platforms.
The observation space for the reinforcement learning agent consists of two components: the normalized laser scan vector and the current map coverage ratio :
This design choice integrates local perception with global state awareness: the laser scan vector captures fine-grained geometric details of the immediate surroundings, furnishing the agent with real-time sensory input for decision-making, while the cumulative map coverage ratio serves as a global indicator of exploration progress, thereby enabling the agent to maintain a coherent understanding of its spatial context.
@@ 3.1.3 Action Space
3.1.4 Neural Networks
We adopt the Proximal Policy Optimization (PPO)[schulman2017proximal] algorithm for policy learning. The observation space is structured as a dictionary comprising -dimensional normalized LiDAR readings and a 1-dimensional map coverage metric, resulting in a ()-dimensional joint input vector. This concatenated observation is processed by a shared backbone network consisting of two fully connected layers, each with 64 neurons and Tanh activation functions. The shared representation is then fed into two separate heads: a policy head that outputs logits over three discrete actions (forward, turn left, turn right), and a value head that estimates the scalar state-value function. All network parameters are jointly optimized in an end-to-end manner via gradient-based updates.
@@ 3.2.1 Lidar Static Anomaly Detection
Robotic exploration can exhibit intentional pausing (e.g., stationary behavior to maximize reward in RL) or motion failure (e.g., collisions or wheel slippage). Undetected motion failures degrade learning efficiency by slowing convergence and reinforcing suboptimal policies. To address this, we propose Lidar Static Anomaly Detection: a lightweight method that identifies motion failures via cosine similarity between consecutive LiDAR scans. Unlike odometry/IMU-based approaches, this method uses raw environmental perception data, maintaining reliability during motor idling or wheel slippage. By operating on normalized scan vectors, it inherently rejects localized environmental changes. Lidar Static Anomaly Detection achieves real-time stagnation detection, with negligible impact on exploration performance.